Coyote Manufacturing bought a delivery truck for $50,000 on July 1, 2017. The registration fee for the truck was $4,000. Coyote also spent $2,000 to paint its logo on the truck. Coyote estimated that the useful life would be four years or 200,000 miles and the residual value at the end of the useful life be $3,000. The mileage to be added to this truck during its useful life are

               2017             2018             2019              2020              2021

       30,000        54,000       52,000          44,000       20,000

Instructions:

  1. Prepare journal entries for the purchase of the truck on July 1. 2017.
  2. Compute depreciation expenses on the truck for the years ending on December 31 of 2017 and 2018     using the following methods;      1) Straight-line method,

2) Activity-based method,

3) Sum-of-the-years'-digit,

4) Double declining method.

3. Coyote sold this truck for $20,000 on January 1, 2019

- Prepare any necessary journal entries for this sale using

1) the sum-of-years-digit method.

2) the double-declining balance method
Answer #1
1 Journal entries:
Date Account titles and explanation Debit Credit
2017
July 1. Delivery truck (50000+4000+2000) 56000
Cash 56000
(Delivery truck purchased)
2 Depreciation expense
Straight-line method:
Depreciation=(Cost-salvage value)/Useful life=(56000-3000)/4=$ 13250
Truck purchased on July 1,2017
Depreciation $
For 2017 Depreciation for 6 months=13250*1/2 6625
For 2018 13250
Activity-based method:
Depreciation per mileage=(Cost-salvage value)/total mileage in the useful life=(56000-3000)/200000=$ 0.265 per lileage
Depreciation expense=Cost per mileage*Miles travelled in the respective years
Depreciation $
For 2017 30000*0.265 7950
For 2018 54000*0.265 14310
Sum-of-the-year's-digit:
Useful life=4 years
Sum of the year's digit=4+3+2+1=10
Depreciation rate for first year=4/10=0.4=40%
Depreciation rate for second year=3/10=0.3=30%
Depreciable cost=Original cost of asset-Salvage value=56000-3000=$ 53000
Depreciation $
For 2017 53000*40%*6/12 (For 6 months) 10600
For 2018 (53000*40%*6/12)+(53000*30%*6/12) 18550
Double declining method:
Depreciation rate=2*(100%/Useful life)=2*(100%/4)=2*25%=50%
Apply depreciation on original cost of asset for the first year
Then apply depreciation on book value
Book value for second year=Original cost-Depreciation for first year=56000-14000=$ 42000
Depreciation $
For 2017 56000*50%*6/12 (For 6 months) 14000
For 2018 42000*50% 21000
3 Date Account titles and explanation Debit Credit
2019
Jan. 1 Cash 20000
Accumulated depreciation (10600+18550) 29150
Loss on sale of delivery truck (Balancing figure) 6850
Delivery truck 56000
(As per sum-of-the-year's-digit method)
Jan. 1 Cash 20000
Accumulated depreciation (14000+21000) 35000
Loss on sale of delivery truck (Balancing figure) 1000
Delivery truck 56000
(As per double-declining balance method)